# Ranking matters: Does the new format select the best teams for the knockout phase in the UEFA Champions League?

### v1

@@ B. Dortmund 15 10 14 12 12 18
The rankings produced by the alternative ranking methods in Section 2 are shown in Table 2. For the generalized row sum, we provide the two extremal results: when the strength of schedule is used only for tie-breaking ($\varepsilon$=0.01) and when the performance of the opponents has the highest possible effect ($\varepsilon=\infty$). The rankings for other values of the parameter $\varepsilon$ can be found in the Appendix, in Table A.1.
Clearly, different methods produce different rankings. With respect to qualification, the clubs are essentially allocated into three sets: ranked in Top 8 (direct qualification for the Round of 16), ranked between 9th and 24th (qualification for the knockout phase play-offs), and ranked between 25th and 36th (elimination). In view of these thresholds, the results can be summarized as follows:
  • Aston Villa falls out of the Top 8 when the strength of schedule has a higher impact. In particular, it only ranks 13th in Keener’s direct ranking method and even only 15th in the GRS with $\varepsilon=\infty$. On the other hand, Real Madrid is ranked 7th according to Keener’s method and GRS ($\varepsilon=\infty$). Hence, it can be argued that Real Madrid should have directly qualified for the Round of 16 at the expense of Aston Villa, even though the latter club has scored one point more.
  • Monaco is clearly the best team among the five with 13 points if the schedule of strength is taken into account, and it can even be ranked higher than some teams with 15 points. Similarly, Benfica could also be ranked higher; it is placed in the Top 8 by Keener’s method.
@@
  • The official ranking, GRS and Keener’s method agree undeniably on the top three teams, which are Liverpool, Barcelona, and Arsenal in this order.
@@
  • All ranking methods agree on the elimination of nine teams (Stuttgart, Shakhtar, Crvena Zvezda, Sturm Graz, Sparta Praha, Girona, Salzburg, S. Bratislava and Young Boys).
In some cases, the ranking methods produce remarkably different results. For example, Atalanta is officially ranked 9th, while it is ranked 18th according to Keener’s direct method. The result of the Colley ranking method is quite similar to the official ranking. As for the GRS ranking, playing only eight rounds in a league with 36 teams seems to be insufficient to obtain a unique ranking. In contrast, in the 2019/20 season of the double round robin European football leagues, where 18–20 teams played 25–29 rounds before the COVID-19 pandemic, the ranking with the generalized row sum method was essentially insensitive to the value of parameter $\varepsilon$ (Csató,, 2021). Indeed, playing a higher number of rounds increases the variance of the number of points and decreases the variance in the strength of opponents, reducing the probability that a team with fewer points scored against stronger opponents is ranked above a team with more points scored against weaker opponents.
@@ 4 Discussion
Possible remedies include (1) using a ranking that reflects the strengths of the teams better than UEFA club coefficients to assign the teams to pots (e.g. a ranking based on Elo ratings as demonstrated by Csató, (2024)), and (2) using the strength of schedule as the primary tie-breaking criterion instead of goal difference (which would boil down to the GRS ranking with, for example, $\epsilon=0.01$). Both recommendations seem to be realistic since the Elo approach has been adopted for the FIFA World Ranking in 2018 (FIFA,, 2018) and the Buchholz method, based on the arithmetic mean of the opponents points’, is the most common tie-breaking rule in Swiss-system tournaments (Freixas,, 2022).
@@ Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.

### v2

@@ B. Dortmund 15 10 14 12 12 12 5
The rankings produced by the alternative ranking methods in Section 2 are shown in Table 2. For the generalized row sum, we provide the two extremal results: when the strength of schedule is used only for tie-breaking ($\varepsilon$=0.01) and when the performance of the opponents has the highest possible effect ($\varepsilon\to\infty$). The rankings for other values of the parameter $\varepsilon$ can be found in the Appendix, in Table A.1.
Clearly, different methods produce different rankings. However, while the direct, Colley, and Generalized Row Sum methods tend to agree, the least squares method give some surprising results. For example, Atleti, with 18 points, is ranked behind Sporting CP that scored only 11 points, and is directly followed by Bologna, with only 6 points. This can be explained by its focus on goal differences as discussed in Section 2.6: the opponents of Atleti and Bologna have a cumulative goal difference of 33-33 and +42+42, respectively. Consequently, least squares indicates that Atleti (Bologna) had an exceptionally (dis)favourable schedule, which is only partially supported by the other methods based on the number of points scored. Therefore, the ranking produced by the least squares will be treated separately in the following discussion.
With respect to qualification, the clubs are essentially allocated into three sets: ranked in the Top 8 (direct qualification for the Round of 16), ranked between 9th and 24th (qualification for the knockout phase play-offs), and ranked between 25th and 36th (elimination). In view of these thresholds, the results can be summarized as follows:
  • Aston Villa falls out of the Top 8 when the strength of schedule has a higher impact. In particular, it is only ranked 13th in the direct ranking method and even only 15th in the GRS with $\varepsilon\to\infty$. On the other hand, Real Madrid is ranked 7th according to the direct method and GRS ($\varepsilon\to\infty$). Hence, it can be argued that Real Madrid should have directly qualified for the Round of 16 at the expense of Aston Villa, even though the latter club has scored one point more.
  • Monaco is clearly the best team among the five with 13 points if the schedule of strength is taken into account, and it can even be ranked higher than some teams with 15 points. Similarly, Benfica could also be ranked higher; it is placed in the Top 8 by the direct method.
@@
Compared to the official ranking, the least squares method removes three teams from the Top 8 and places three teams with 15 points directly in this set. Furthermore, Feyenoord would be eliminated instead of Bologna, which would be difficult to explain for the stakeholders since the former team has 13 points rather than 6. Last but not least, even Leipzig would qualify for the play-offs despite its sole win against Sporting CP.
@@
  • All ranking methods agree on the top three teams, which are Liverpool, Barcelona, and Arsenal.
@@
  • All ranking methods agree on the elimination of nine teams (Stuttgart, Shakhtar, Crvena Zvezda, Sturm Graz, Sparta Praha, Girona, Salzburg, S. Bratislava, and Young Boys).
In some cases, the ranking methods produce remarkably different results even if the least squares method is disregarded. For example, Atalanta is officially ranked 9th, while it is ranked 18th according to the direct method. The result of the Colley ranking method is quite similar to the official ranking, which can be explained by its strong relation to the GRS as presented in Section 2.5.1.
As for the GRS ranking, playing only eight rounds in a league with 36 teams seems to be insufficient to obtain a unique ranking. In contrast, in the 2019/20 season of the double round robin European football leagues, where 18–20 teams played 25–29 rounds before the COVID-19 pandemic, the ranking with the generalized row sum method was essentially insensitive to the value of parameter $\varepsilon$ (Csató,, 2021). Indeed, playing a higher number of rounds increases the variance of the number of points and decreases the variance in the strength of opponents, reducing the probability that a team with fewer points scored against stronger opponents is ranked above a team with more points scored against weaker opponents.
@@ 4 Discussion
Possible remedies include (1) using a ranking that reflects the strengths of the teams better than UEFA club coefficients to assign the teams to pots (e.g. a ranking based on Elo ratings as demonstrated by Csató, (2024)), and (2) using the strength of schedule as the primary tie-breaking criterion instead of goal difference (which would boil down to the GRS ranking with, for example, $\varepsilon=0.01$). Both recommendations seem to be realistic since the Elo approach has been adopted for the FIFA World Ranking in 2018 (FIFA,, 2018) and the Buchholz method, based on the arithmetic mean of the points scored by opponents, is the most common tie-breaking rule in Swiss-system tournaments (Freixas,, 2022).
@@ Acknowledgements
Four reviewers (Julien Guyon, Mehmet S. Ismail, Michael A. Lapré, Fernando Leiva Bertrán), and Kees Pippel provided valuable remarks on an earlier draft.
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838, by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences, and by the Research Foundation - Flanders (FWO) under grant no. S005521N.